You are given the following information: Stockholders' equity as reported on the firm’s balance sheet = $2.5 billion, price/earnings ratio = 10.5, common shares outstanding = 180 million, and market/book ratio = 2.4. The firm's market value of total debt is $7 billion, the firm has cash and equivalents totaling $320 million, and the firm's EBITDA equals $1 billion. What is the price of a share of the company's common stock? Do not round intermediate calculations. Round your answer to the nearest cent.

$  

What is the firm's EV/EBITDA? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Given,

Book value of Shareholder's equity = $2.5 billion

P/E = 10.5

Common shares outstanding = 180 million

Market to book value of equity = 2.4

So, Market to book value of equity = Market Value per share /Book value per share

Book value per share = Book value of Shareholder's equity/ Common shares outstanding

BVPS = 2500/180 = $13.89

So, Market value = 2.4 * 13.89 = $33.33

the price of a share of the company's common stock = $33.33

Given EBITDA = $1 Billion

market value of total debt = $7 billion,

cash and equivalents = $320 million

Enterprise Value = Market value of Debt + Market value of equity - cash

SO, EV = 7000 + 2500 - 320 = $9180 million

EV/EBITDA = 9180/1000 = 9.18x